Find the particular solution of nonlinear differential equation 

𝑑𝑦

𝑑𝑥

=

𝑦

2−1

𝑥

, 𝑦(1) = 2.



1
Expert's answer
2021-05-07T13:09:28-0400

"\displaystyle
y' = \frac{y^2 - 1}{x}\\

\frac{\mathrm{d}y}{y^2 - 1} = \frac{\mathrm{d}x}{x}\\

\int\frac{\mathrm{d}y}{y^2 - 1} = \int\frac{\mathrm{d}x}{x}\\

\ln\left(\frac{y - 1}{y + 1}\right) = 2\ln{x} + C\\

\frac{y - 1}{y + 1} = Ax^2\\

y - 1 = Ax^2(y + 1)\\

y(1 - Ax^2) = Ax^2 + 1\\

y = \frac{Ax^2 + 1}{1 - Ax^2}\\

y(1) = \frac{1 + A}{1 - A} = 2\\

1 + A = 2 - 2A\\

3A = 1, A = \frac{1}{3}\\


\therefore y = \frac{x^2 + 3}{3 - x^2}"